. The stone with a mass of 200g is thrown up with an initial velocity V0 = 6 m / s. á) What happens to the Impulse during climbing, increases or decreases? Why? b) Calculate the change in momentum as the stone reaches its greatest height.


1
Expert's answer
2021-01-31T18:58:45-0500

a) Impulse decreases. It is from


"mv_f-mv_i=-mgt"

b)

"\Delta p=-mv=-(0.2)(6)=-1.2\frac{kg\cdot m}{s}"